OpenStudy (hoybaby):

determine whether the sequence is converging or diverging. will post question, click this please

OpenStudy (hoybaby):

\[a _{n}= \frac{ e ^{n}+e ^{-n} }{ e ^{2n}-3 }\]

OpenStudy (hoybaby):

having a tough time determining

OpenStudy (vuriffy):

I believe that's converging. One second.

OpenStudy (hoybaby):

towards what exactly?

OpenStudy (hoybaby):

is it diveriging to 1? i think it is

OpenStudy (hoybaby):

because the answer format is the equation i listed but with this = (blank)

OpenStudy (hoybaby):

i tried putting that answer in the webassign but showed i was wrong

OpenStudy (vuriffy):

I'm not completely sure, it's been a few years, so I apologize that I cannot help fully.

OpenStudy (vuriffy):

It may actually be converging.

OpenStudy (dumbcow):

if you take limit as n-> infinity, the limit = 0. Series converges to 0 \[\lim_{n \rightarrow \infty} \frac{e^n + \frac{1}{e^n}}{e^{2n} -3} = \lim_{n \rightarrow \infty} \frac{1}{e^n} (\frac{e^{2n}+1}{e^{2n}-3}) = 0*(1) = 0\]